A vacuum is a space in which there is no substance. There is not quite a complete vacuum in space either. The average pressure in space is of the order of 10 –15 Pa. How many gas molecules are in space in one cubic meter when the temperature of space is 3 K?
1
Expert's answer
2020-05-05T18:41:22-0400
"pV=\frac{N}{N_A}RT"

"(10^{-15})(1)=\frac{N}{6\cdot 10^{23}}(8.31)(3)"

"N=2.4\cdot 10^{7}"